OpenStudy (anonymous):

The sum of twice the larger number and the smaller number is 43. Using "L" to represent the larger number, translate and simplify the following expression in terms of L: Three times the smaller number less than the sum of the larger number and four.

OpenStudy (anonymous):

The first part tells you how L and s ('s' for 'smaller') relate. 'Twice the larger and the smaller is 43' means\[2L+s=43\]The next part says, 'Three times the smaller less than the sum of the larger and four' which means\[L+4-3s\]You're told to express this last one in terms of L only (i.e. get rid of s). You use the first equation to do that. \[s=43-2L\]and so you have\[L+4-3s\]\[=L+4-3(43-2L)\]\[=L+4-129+6L\]\[=7L-125\]

OpenStudy (anonymous):

YES! s=43-2L then replace the s w/that!

OpenStudy (anonymous):

Yes.